The volume of a balloon is 3.02 L at 22.7°C.  The balloon is heated to 43.6°C.  Calculate the new volume of the balloon.

Question 7 options:

A)

1.57 L

B)

3.02 L

C)

2.82 L

D)

3.23 L

E)

5.80 L
Answer #1

Answer :

Correct option is (D) 3.23 L

Concept : Charle's law

According to Charle's law , Volume increases on increasing in temperature keeping other parameters like pressure and moles of gas constant.

Formule for relating 2 Volume with 2 temperature

V1 / T1 = V2 / T2

Given ,

Initial volume (V1) = 3.02 L

Initial temperature (T1) = 22.7°C = 295.7 K

( Since , 0°C = 273 K)

T2 ( final temperature) = 43.6 °C = 316.6 K

So,

V2 = ( V1 / T1) * T2

= ( 3.02 L / 295.7 K) * 316.6 K

= 3.23 L

So, new Volume of balloon = 3.23 L